### v1

@@ 5 Numerical Experiments
In Tables 1 and 2, we report the Lipschitz constant of the gradient of $f$ (denoted as Lips), the number of outer iterations (O-IT), the number of inner iterations (I-IT), the number of linesearch iterations (LS-IT), and the total running time in seconds (Time). The results indicate that, in terms of CPU times, the IPG-ELS method outperforms the other two methods. This efficiency can be attributed to two main factors: (i) the PG-ELS method requires a significantly higher number of inner iterations to solve the proximal subproblem ”exactly”, and (ii) the IPG-FixStep method employs a small stepsize of $1/L$ in the gradient step.
@@
Colon Tumor 41.58 IPG-ELS 1.1056 101 195 318 28.88
PG-ELS 1.1056 104 751 324 36.19
IPG-FixStep 1.1056 1450 1451 - 171.72
Colon Tumor 665.32 IPG-ELS 2.3647 101 101 823 32.74
PG-ELS 2.3623 128 455 1038 44.02
IPG-FixStep 2.3647 1103 1104 - 97.72
Colon Tumor 5133.69 IPG-ELS 5.8989 101 101 1134 110.86
PG-ELS 5.8832 119 328 1324 116.26
IPG-FixStep 5.8981 586 587 - 138.92
Heart Disease 77.12 IPG-ELS 0.1732 101 178 539 1.46
PG-ELS 0.1732 119 876 603 4.28
IPG-FixStep 0.1732 487 488 - 3.52
Heart Disease 1233.99 IPG-ELS 0.3129 101 101 903 0.85
@@
IPG-FixStep 0.3520 2001 2001 - 11.85
Heart Disease 9521.59 IPG-ELS 0.4995 101 101 1222 0.99
PG-ELS 0.4992 227 622 2690 5.46
IPG-FixStep 0.7242 2001 2001 - 43.92
CNS 41.95 IPG-ELS 0.9519 101 182 341 364.44
PG-ELS 0.9518 119 714 396 501.59
IPG-FixStep 0.9519 1217 1218 - 1675.05
CNS 671.17 IPG-ELS 2.1153 101 101 768 544.91
PG-ELS 2.1119 148 725 1110 878.39
IPG-FixStep 2.2193 2001 2001 - 2749.17
CNS 5178.80 IPG-ELS 6.0665 101 101 1130 1089.70
PG-ELS 6.0613 102 343 1138 1169.25
IPG-FixStep 6.0655 775 776 - 1677.93
Table 1: Performance of the IPG-ELS, PG-ELS and IPG-FixStep methods on $3$ data sets.
@@
Lung cancer 52.58 IPG-ELS 0.8985 101 179 443 422.07
PG-ELS 0.8984 105 837 457 521.46
IPG-FixStep 0.8985 678 679 - 962.80
Lung cancer 841.22 IPG-ELS 2.7632 101 101 845 591.92
PG-ELS 2.7623 131 768 1085 851.75
IPG-FixStep 2.7631 588 589 - 838.30
Lung cancer 2658.70 IPG-ELS 3.6391 101 101 992 696.61
PG-ELS 3.6378 161 740 1574 1173.05
IPG-FixStep 3.8819 2001 2001 - 2969.58
Secom 45.78 IPG-ELS 0.6438 101 175 373 66.64
PG-ELS 0.6438 99 9247 360 694.81
IPG-FixStep 0.6438 857 858 - 225.29
Secom 732.50 IPG-ELS 0.8587 101 101 779 82.81
PG-ELS 0.8586 102 4931 795 423.80
IPG-FixStep 0.8586 1662 1663 - 440.13
Secom 5652.06 IPG-ELS 1.6981 101 101 1138 104.54
PG-ELS 1.6899 125 1346 1388 215.08
IPG-FixStep 1.6977 801 802 - 211.93
Cina0 68.39 IPG-ELS 0.7972 101 245 487 3041.60
PG-ELS 0.7972 104 1490 483 5005.84
IPG-FixStep 0.7972 608 790 - 5530.44
Cina0 527.70 IPG-ELS 1.2817 101 250 838 3925.72
PG-ELS 1.2817 94 1789 693 5712.33
IPG-FixStep 1.2817 1010 1011 - 8741.06
Cina0 8443.20 IPG-ELS 3.6531 101 104 1126 3791.25
PG-ELS 3.6530 284 3083 3061 15338.18
IPG-FixStep 3.8493 2001 2001 - 17104.23
Table 2: Performance of the IPG-ELS, PG-ELS and IPG-FixStep methods on $3$ data sets.
@@ Acknowledgements.
YBC was partially supported by the NSF Grant DMS-2307328, and by an internal grant from NIU. MLNG was partially supported by CNPq Grants 405349/2021-1 and 304133/2021-3. JGM was partially supported by CNPq Grant 312223/2022-6.

### v2

@@ 6 Numerical Experiments
In Tables 1 and 2, we report the Lipschitz constant of the gradient of $f$ (denoted as Lips), the number of outer iterations (O-IT), the number of inner iterations (I-IT), the number of linesearch iterations (LS-IT), and the total running time in seconds (Time). The results indicate that, in terms of CPU times, the IPG-ELS method outperforms the other three methods. This efficiency can be attributed to two main factors: (i) the PG-ELS and Tseng-MFBS methods require significantly more inner iterations to solve the proximal subproblem ”exactly”, and (ii) the IPG-FixStep and Tseng-MFBS methods employ small stepsizes of $1/L$ and $0.9/L$, respectively, in the gradient step.
@@
Colon Tumor 41.58 IPG-ELS 1.1056 101 195 318 25.80
PG-ELS 1.1056 104 751 324 32.82
IPG-FixStep 1.1056 1450 1451 - 137.78
Tseng-MFBS 1.1056 1615 4845 - 193.92
Colon Tumor 665.32 IPG-ELS 2.3647 101 101 823 31.02
PG-ELS 2.3623 128 455 1038 43.40
IPG-FixStep 2.3647 1103 1104 - 94.51
Tseng-MFBS 2.3645 1227 2454 - 116.80
Colon Tumor 5133.69 IPG-ELS 5.8989 101 101 1134 49.30
PG-ELS 5.8832 119 328 1324 63.73
IPG-FixStep 5.8981 586 587 - 66.26
Tseng-MFBS 5.8980 652 1304 - 79.46
Heart Disease 77.12 IPG-ELS 0.1732 101 178 539 1.25
PG-ELS 0.1732 119 876 603 3.68
IPG-FixStep 0.1732 487 488 - 2.92
Tseng-MFBS 0.1732 541 1082 - 4.79
Heart Disease 1233.99 IPG-ELS 0.3129 101 101 903 0.86
@@
IPG-FixStep 0.3520 2001 2001 - 11.87
Tseng-MFBS 0.3577 2001 4002 - 17.56
Heart Disease 9521.56 IPG-ELS 0.4995 101 101 1222 1.01
PG-ELS 0.4992 227 622 2690 3.80
IPG-FixStep 0.7242 2001 2001 - 14.85
Tseng-MFBS 0.7533 2001 4002 - 20.74
CNS 41.95 IPG-ELS 0.9519 101 182 341 397.09
PG-ELS 0.9518 119 714 396 531.26
IPG-FixStep 0.9519 1217 1218 - 1762.69
Tseng-MFBS 0.9519 1354 3091 - 2263.33
CNS 671.17 IPG-ELS 2.1153 101 101 768 554.27
PG-ELS 2.1119 148 725 1110 885.04
IPG-FixStep 2.2193 2001 2001 - 2752.47
Tseng-MFBS 2.2439 2001 4002 - 2828.31
CNS 5178.78 IPG-ELS 6.0665 101 101 1130 854.13
PG-ELS 6.0613 102 343 1138 896.40
IPG-FixStep 6.0655 775 776 - 1240.28
Tseng-MFBS 6.0653 862 1724 - 1430.63
Table 1: Performance of the IPG-ELS, PG-ELS, IPG-FixStep and Tseng-MFBS methods on $3$ data sets.
@@
Lung cancer 52.58 IPG-ELS 0.8985 101 179 443 539.37
PG-ELS 0.8984 105 837 457 646.59
IPG-FixStep 0.8985 678 679 - 1216.98
Tseng-MFBS 0.8985 755 2265 - 1487.95
Lung cancer 841.23 IPG-ELS 2.7632 101 101 845 601.68
PG-ELS 2.7623 131 768 1085 875.26
IPG-FixStep 2.7631 588 589 - 850.73
Tseng-MFBS 2.7631 654 1308 - 978.97
Lung cancer 2658.70 IPG-ELS 3.6391 101 101 992 711.25
PG-ELS 3.6378 161 740 1574 1398.46
IPG-FixStep 3.8819 2001 2001 - 3437.99
Tseng-MFBS 3.9389 2001 4002 - 3124.96
Secom 45.78 IPG-ELS 0.6438 101 175 373 99.15
PG-ELS 0.6438 99 9247 360 822.17
IPG-FixStep 0.6438 857 858 - 304.24
Tseng-MFBS 0.6438 952 34344 - 5216.87
Secom 732.51 IPG-ELS 0.8587 101 101 779 86.11
PG-ELS 0.8586 102 4931 795 431.53
IPG-FixStep 0.8586 1662 1663 - 454.59
Tseng-MFBS 0.8587 1847 3969 - 637.36
Secom 5652.07 IPG-ELS 1.6981 101 101 1138 108.07
PG-ELS 1.6899 125 1346 1388 220.10
IPG-FixStep 1.6977 801 802 - 218.15
Tseng-MFBS 1.6977 891 1782 - 296.50
Cina0 68.39 IPG-ELS 0.7972 101 245 487 3168.99
PG-ELS 0.7972 104 1490 483 5925.88
IPG-FixStep 0.7972 609 610 - 6451.18
Tseng-MFBS 0.7972 677 1354 - 8000.69
Cina0 527.70 IPG-ELS 1.2817 101 250 838 4567.13
PG-ELS 1.2817 94 1789 693 5704.39
IPG-FixStep 1.2817 1010 1011 - 8735.91
Tseng-MFBS 1.2817 1122 2244 - 11564.76
Cina0 8443.20 IPG-ELS 3.6531 101 104 1126 3968.04
PG-ELS 3.6530 284 3083 3061 15889.08
IPG-FixStep 3.8493 2001 2001 - 17796.07
Tseng-MFBS 3.8657 2001 4002 - 20681.39
Table 2: Performance of the IPG-ELS, PG-ELS, IPG-FixStep and Tseng-MFBS methods on $3$ data sets.
@@ 7 Conclusions
Declarations
Funding
YBC
was partially supported by the NSF Grant DMS-2307328, and by an internal grant from NIU. MLNG was partially supported by CNPq Grants 405349/2021-1 and 304133/2021-3. JGM was partially supported by CNPq Grant 312223/2022-6.
Data availability statement
The codes supporting the numerical experiments are freely available in the homepage:
https://maxlng.ime.ufg.br/p/17888-publications.
Conflicts of interest
The authors declare that they have no conflict of interest.
@@ Acknowledgements.
The authors are grateful to the anonymous referees for their valuable comments and suggestions which helped to improve the quality of the paper.